OpenStudy (anonymous):

Convert the following equation from rectangular to polar form: y = x^2

OpenStudy (anonymous):

y=rsin(theta) x=xcos(theta) So: rsin(theta)=r^2cos^2(theta)

OpenStudy (anonymous):

r=cot(theta)csc(theta) r =tan(theta)sec(theta) r =tan(theta) r = cot(theta) those are the answer choices

OpenStudy (a_clan):

\[r=\sqrt{x ^{2}+y ^{2}} \] \[\tan \theta=y/x\]

OpenStudy (anonymous):

You want to divide out the common factor of r giving: sin(theta)=rcos^2(Theta) Divide over the cos^2(theta) sin(theta)/cos(theta)*(1/cos(theta))=r so r=tan(theta)sec(theta)

OpenStudy (a_clan):

So \[\tan \theta=x ^{2}/x=x\]

OpenStudy (a_clan):

now \[r=\sqrt{x ^{2}+x ^{4}} = x \sqrt{1+x ^{2}}\]

OpenStudy (a_clan):

or \[r=\tan \theta \sqrt{1+\tan ^{2}\theta}\]

OpenStudy (a_clan):

now I suppose it can be simplified to an answer

OpenStudy (anonymous):

Of: \[r=\tan(\theta)\sec(\theta)\]